bob has $2.35 in nickles & dimes. He has 33 coins in all. Find the number of nickles and dimes. solve.
Answer by flame8855(424)

x be the number of dimes
y be the number of nickels
0.1x+0.05y=2.35
x+y = 33
solve the two equations
multiple the first by 100
10x+ 5y = 235
x+y= 33
x= 14
y= 19
14 dimes & 19 nickels